OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 15 meters and legs that measure 9 meters and 12 meters. What is the surface area of this prism if the height of the prism is 11 meters?

Directrix (directrix):

Assuming that the prism is a right triangular base prism, its surface area is the sum of its lateral area and the area of its two triangular bases. SA = ph + 2 B where p is the perimeter of the base and B is the area of each base. SA = ( 9 + 12 + 15) * 11 + 2 (1/2) (9)(12) SA = 36 * 11 + 2 (108) = 612 square meters

OpenStudy (lexx):

its not 612!